Hello. I am having problems with this 1 question.
How many mL of water should be added to 1 Gallon of 70% isopropyl alcohol to prepare a 30% solution for soaking sponges?
I would appreciate your help,,, Thank You !!!!

Answer by Alan3354(69443)

How many mL of water should be added to 1 Gallon of 70% isopropyl alcohol to prepare a 30% solution for soaking sponges?
-------------
1*70 + w*0 = (1+w)*30
70 = 30 + 30w
w = 4/3 gallons of water